# Counting primes with a given primitive root, uniformly

### v1

@@ 1. Introduction
for all $z\geq 1$ and $a\in\mathbb{N}$. Since an application of [11, Eq. (3.18)] yields
$\sum_{m\leq(x/\ell)^{\beta}/V}\frac{\mu(m)^{2}}{\varphi(m)}>\log\left((x/\ell)^{\beta}/V\right)=\left(\frac{1}{2}+O\left(\frac{1}{\log\log x}\right)\right)\log(x/\ell),$
@@
Inserting this and (23) in (22) yields
@@
Combining the above with (21), we find that the count of $p\leq x$ corresponding to a given $\ell\in J_{3}$ is at most
@@
By Mertens’ theorem and partial summation, we have
$\sum_{\ell\in J_{3}}\frac{1}{\ell\log(x/\ell)}$ $=\int_{t\in J_{3}}\frac{1}{\log(x/t)}\,\mathrm{d}\left(\sum_{\ell\leq t}\frac{1}{\ell}\right)$
$=\int_{t\in J_{3}}\frac{\mathrm{d}t}{t(\log t)\log(x/t)}+\int_{t\in J_{3}}\frac{1}{\log(x/t)}\,\mathrm{d}\left(O\left(\frac{1}{\log t}\right)\right)$
@@
Finally, it remains to estimate the number of primes $p\in\mathcal{S}$ failing the $\ell$-test for some $\ell\in J_{4}$. For each such $p$, the order of $g\pmod{p}$ is smaller than $x^{1-\alpha}$. Thus, $p\mid(g^{m}-1)$ for some positive integer $m\leq x^{1-\alpha}$. The number of distinct prime factors of $g^{m}-1$ is $O(m\log|g|)$. Hence, the number of primes $p\in\mathcal{S}$ failing the $\ell$-test for some $\ell\in J_{4}$ is at most
@@
Since $\alpha\in(10/19,1)$, we have $12-2\alpha+1/B<$ 1. Thus, $x^{2-2\alpha}\log|g|$ is of smaller order than the main term in (20).
@@
Since $\tilde{A}_{1}(g)\geq\tilde{A}_{1}(21)=4/5$ for $g>1$, the proof of Theorem 1.2 shows that the exponent $B=19$ can be improved to $16$ if we focus merely on positive $g\in\mathcal{G}$. Besides, if the square factor of $g$ has size $o(|g|)$ or $g_{1}\not\equiv 1\pmod{4}$, we have $\tilde{A}_{1}(g)=1+o(1)$ for $g$ with $|g|$ sufficiently large. Consequently, our proof of Theorem 1.2 yields $p_{g}\ll\log^{13}(2|g|)$ for these $g\in\mathcal{G}$. In particular, this inequality holds for all squarefree $g\in\mathcal{G}$.
@@
The following lemma is due to Vaughan (see Theorem 4.1 in [17]).
@@
For a certain constant $\alpha>0$, we have
$\sum_{2<p\leq y}\frac{\varphi(p-1)}{p-\varphi(p-1)}=(\alpha+o(1))\frac{y}{\log{y}},\qquad\text{as $y\to\infty$}.$
Put $L=\log{x}/\log\log{x}$. Let $\delta_{p}$ be defined as in (6), and put $M_{p}=\prod_{r\leq p}r$. Then $p_{g}=p$ precisely when $g$ belongs to one of $\delta_{p}M_{p}$ residue classes modulo $M_{p}$. Since $M_{p}\ll 3^{p}$,
#$\#\{g:|g|\leq x:p_{g}=p\}=2\delta_{p}x+O(3^{p}).$
As #$\#[-x,x]\setminus\mathcal{G}\ll x^{1/2}$, it follows that
$\sum_{\begin{subarray}{c}g\in\mathcal{G}\\ |g|\leq x\\ p_{g}\leq L\end{subarray}}p_{g}=\sum_{p\leq L}p\sum_{\begin{subarray}{c}g\in\mathcal{G}\\ |g|\leq x\\ p_{g}=p\end{subarray}}1$ $=2x\sum_{p\leq L}p\delta_{p}+O\bigg{(}\sum_{p\leq L}p(3^{p}+x^{1/2})\bigg{)}$
$=2x\sum_{p\leq L}p\delta_{p}+O(x^{1/2}L^{2}).$ (24)
We now extend the sum on $p$ to infinity, using Lemma 4.1 to estimate the resulting error. Observe that
$\delta_{p}<\prod_{r<p}\left(1-\frac{\varphi(r-1)}{r}\right)=\prod_{r<p}\left(1+\frac{\varphi(r-1)}{r-\varphi(r-1)}\right)^{-1}.$
If $r>2$, then $r-1$ is even, and $\varphi(r-1)\leq\frac{r-1}{2}$. Hence, $r-\varphi(r-1)>\varphi(r-1)$, and the ratio $\frac{\varphi(r-1)}{r-\varphi(r-1)}<1$. Using the inequality $1+u\geq\exp(u/2)$ valid when $0\leq u\leq 1$, we conclude from Lemma 4.1 that for all sufficiently large $p$,
@@
$\prod_{r<p}\left(1+\frac{\varphi(r-1)}{r-\varphi(r-1)}\right)\geq\exp\left(\frac{1}{2}\sum_{2<r<p}\frac{\varphi(r-1)}{r-\varphi(r-1)}\right)\geq\exp(cp/\log{p}),$
for $c:=\frac{1}{3}\alpha$. As a consequence, $\delta_{p}\ll\exp(-cp/\log{p})$ for all primes $p$, and
@@
Referring back to (24), we deduce that
$\sum_{\begin{subarray}{c}g\in\mathcal{G}\\ |g|\leq x\\ p_{g}\leq L\end{subarray}}p_{g}=2x\sum_{p}p\delta_{p}+O(x\exp(-(\log{x})^{1+o(1)})).$
Next, we bound the sum of the $p_{g}$ taken over $g\in\mathcal{G}$, $|g|\leq x$, having $p_{g}>L$. If $p_{g}>L$, then $g$ belongs to one of $\gamma M$ residue classes mod $M$, where
$M:=\prod_{r\leq L}r,\quad\text{and}\quad\gamma:=\prod_{r\leq L}\left(1-\frac{\varphi(r-1)}{r}\right).$
The number of such $g$ with $|g|\leq x$ is $\ll\gamma(x+M)\ll\gamma x$, noting that $M\leq 3^{L}=x^{o(1)}$. Moreover, essentially the same work used to estimate $\gamma_{p}$ shows that $\gamma\leq\exp(-cL/\log{L})$. (All of this is being claimed for large enough values of $x$.) So by Theorem 1.2,
$\sum_{\begin{subarray}{c}g\in\mathcal{G}\\ |g|\leq x\\ p_{g}>L\end{subarray}}p_{g}$ $\leq(\max_{\begin{subarray}{c}g\in\mathcal{G}\\ |g|\leq x\end{subarray}}p_{g})\sum_{\begin{subarray}{c}g\in\mathcal{G}\\ |g|\leq x\\ p_{g}>L\end{subarray}}1\ll(\log{x})^{16}(x\exp(-cL/\log{L}))\ll x\exp(-(\log{x})^{1+o(1)}).$
Putting together the pieces,
$\sum_{\begin{subarray}{c}g\in\mathcal{G}\\ |g|\leq x\end{subarray}}p_{g}=2x\sum_{p}p\delta_{p}+O(x/\exp((\log{x})^{1+o(1)})).$
Corollary 1.3 follows; in fact, the ratio appearing on the left in (7) is $\sum_{p}p\delta_{p}+O(\exp(-(\log{x})^{1+o(1)}))$.
@@
Our main tool for this proof will be Montgomery’s “arithmetic large sieve” inequality introduced in Section 3. Using Montgomery’s sieve, Vaughan showed [17] that $p_{g}\leq N^{1/2}$ for all $g\in[M+1,M+N]$, apart from $O(N^{1/2}(\log{N})^{1-\alpha})$ exceptions, where $\alpha$ is the constant of Lemma 4.1. Earlier Gallagher [5] had shown such a result with $1$ in place of $1-\alpha$. The next proposition implies that $N^{1/2}$ can be replaced by a large power of $\log{N}$, if one is willing to slightly inflate the exponent $1/2$ on $N$ in the size of the exceptional set.
@@
We may assume $N$ is sufficiently large. We apply the arithmetic large sieve with $Q=N^{1/2}$, taking $\nu(p)=\varphi(p-1)$ for $p\leq Y$, and $\nu(p)=0$ for $Y<p\leq Q$. It suffices to show that with these choices of parameters, the denominator
$J=\sum_{n\leq N^{1/2}}\mu^{2}(n)\prod_{p\mid n}\frac{\varphi(p-1)}{p-\varphi(p-1)}.$ (25)
@@
$J\geq N^{1/2}\exp\left(O\left(\log{N}\frac{\log\log\log{N}}{\log\log{N}}\right)\right)\cdot\exp(-u\log{u}).$ (26)
Let $R$ be the number of primes $p\in[\frac{1}{2}Y,Y]$ for which the smallest prime factor of $\frac{p-1}{2}$ exceeds $Y^{1/5}$. By the linear sieve and the Bombieri–Vinogradov theorem,
$R>Y/\log^{3}{Y}.$
(We need for the application of Bombieri–Vinogradov that $\frac{1}{5}<\frac{1}{4}$.) For each such $p$, the ratio $\frac{\varphi(p-1)}{p-1}=\frac{1}{2}\prod_{\ell\mid p-1,~{}\ell>2}(1-1/\ell)>\frac{1}{2}(1-y^{-1/5})^{4}>2/5$ (say). Hence, $\frac{\varphi(p-1)}{p}>\frac{1}{3}$, and $\frac{\varphi(p-1)}{p-\varphi(p-1)}>\frac{1}{2}$. Let $u_{0}=\lfloor\log(N^{1/2})/\log{Y}\rfloor$ (so that $u_{0}=\lfloor u\rfloor$, with $u$ as in the proposition). By considering the contribution to the right-hand side of (25) from products of $u_{0}$ distinct primes $p$ of the above kind, we see that $J\geq 2^{-u_{0}}\binom{R}{u_{0}}$. Now $R>Y/(\log{Y})^{3}>(\log{N})^{3/2}>u_{0}$. Since $\binom{n}{k}\geq(n/k)^{k}$ for each pair of integers $n,k$ with $n\geq k>0$, we conclude that
@@
The assumed bounds on $Y$ ensure that $Y^{-1}(2(\log{Y})^{3})^{-u}=\exp\left(O\left(\log{N}\frac{\log\log\log{N}}{\log\log{N}}\right)\right)$. Our desired lower estimate (26) follows by combining the last two displays. ∎
@@
Fix $K\geq 2$ with $\eta+\frac{1}{2}(1+1/K)<1$. We first consider the contribution of $g\in\mathcal{G}$, $|g|\leq x$, having $p_{g}\leq\log^{K}{(3x)}$. Note that the corresponding summand $\min\{p_{g},x^{\frac{1}{2}-\varepsilon}\}=p_{g}$ for these values of $g$ (once $x$ exceeds a certain constant depending only on $K$).
In the course of proving Corollary 1.3, we showed that with $L=\log{x}/\log\log{x}$,
@@
Furthermore, the count of $g\in\mathcal{G}$, $|g|\leq x$ with $p_{g}>L$ is $O(x\exp(-cL/\log{L}))$. Hence,
@@
Therefore, the theorem will be proved if is shown that
$\sum_{\begin{subarray}{c}g\in\mathcal{G},~{}|g|\leq x\\ p_{g}>\log^{K}(3x)\end{subarray}}\min\{p_{g},x^{\frac{1}{2}-\varepsilon}\}=o(x),$
as $x\to\infty$. For this we apply Proposition 5.1. Choose $M$ and $N$ with $M+1=-\lfloor x\rfloor$ and $M+N=\lfloor x\rfloor$; then $[M+1,M+N]$ is the set of all integers $g$ with $|g|\leq x$, and $N=2\lfloor x\rfloor+1<3x$. Thus, if $p_{g}>\log^{K}{(3x)}$, then $p_{g}>\log^{K}{N}$. By Proposition 5.1, the number of such $g$, $|g|\leq x$, is at most $x^{\frac{1}{2}(1+1/K)+o(1)}$. It follows that the sum appearing in the last display is bounded above by $x^{\eta}\cdot x^{\frac{1}{2}(1+1/K)+o(1)}$, which is $o(x)$ by our choice of $K$. ∎

### v2

@@ 2. Notation
for all $z\geq 1$ and $a\in\mathbb{N}$. Since an application of [13, eq. (3.18)] yields
$\sum_{m\leq(x/\ell)^{\beta}/V}\frac{\mu(m)^{2}}{\varphi(m)}>\log\frac{(x/\ell)^{\beta}}{V}=\left(\frac{1}{2}+O\left(\frac{1}{\log\log x}\right)\right)\log(x/\ell),$
@@
Inserting this estimate and (4.5) in (4.4) yields
@@
Combining the above with (4.3), we find that the count of $p\leq x$ corresponding to a given $\ell\in J_{3}$ is at most
@@
By Mertens’ theorem and partial summation [13, eq. (A.4), p. 488], we have
$\sum_{\ell\in J_{3}}\frac{1}{\ell\log(x/\ell)}$ $=\int_{J_{3}}\frac{1}{\log(x/t)}\,\mathrm{d}\left(\sum_{\ell\leq t}\frac{1}{\ell}\right)$
$=\int_{J_{3}}\frac{\mathrm{d}t}{t(\log t)\log(x/t)}+\int_{J_{3}}\frac{1}{\log(x/t)}\,\mathrm{d}\left(O\left(\frac{1}{\log t}\right)\right)$
@@
Finally, it remains to estimate the number of primes $p\in\mathcal{S}$ failing the $\ell$-test for some $\ell\in J_{4}$. For each such $p$, the order of $g\bmod{p}$ is smaller than $x^{1-\alpha}$. Thus, $p\mid(g^{m}-1)$ for some positive integer $m\leq x^{1-\alpha}$. The number of distinct prime factors of $g^{m}-1$ is $O(m\log|g|)$. Hence, the number of primes $p\in\mathcal{S}$ failing the $\ell$-test for some $\ell\in J_{4}$ is at most
@@
Since $\alpha\in(10/19,1)$, we have $2-2\alpha+1/B<1$. Thus, $x^{2-2\alpha}\log|g|$ is of smaller order than the main term in (4.2).
@@
Since $\tilde{A}_{1}(g)\geq\tilde{A}_{1}(21)=4/5$ for $g>1$, the proof of Theorem 1.2 shows that the exponent $B=19$ can be improved to $16$ if we focus merely on positive $g\in\mathcal{G}$. Besides, if we write $g=g_{1}m^{2}$ with $g_{1}\in\mathbb{Z}$ squarefree and $m\in\mathbb{N}$, then $\tilde{A}_{1}(g)=1+o(1)$ whenever $|g|$ is sufficiently large, provided that $m^{2}=o(|g|)$ or $g_{1}\not\equiv 1\pmod{4}$. Consequently, our proof of Theorem 1.2 yields $p_{g}\ll\log^{13}(2|g|)$ for these $g\in\mathcal{G}$. In particular, this inequality holds for all squarefree $g\in\mathcal{G}$.
@@
We remind the reader that $r$ is always to be understood as representing a prime number. We let $r_{1}=2,r_{2}=3,r_{3}=5,\dots$ denote the sequence of primes in the usual increasing order.
@@
For a certain constant $\varrho>1$, we have
@@
$\prod_{r\leq r_{k}}\left(1-\frac{\varphi(r-1)}{r}\right)=\varrho^{-(1+o(1))k}\quad\text{as}\quad k\to\infty.$
Lemma 5.1 and the prime number theorem together imply that
$\prod_{r\leq y}\left(1-\frac{\varphi(r-1)}{r}\right)=\exp(-(1+o(1))(\log{\varrho})y/\log{y}),$ (5.1)
as $y\to\infty$. We make repeated use below of this form of Lemma 5.1.
Proof of Lemma 5.1.
We will prove the lemma for a constant $\varrho$ constructed in terms of the moments of $\frac{\varphi(r-1)}{r-1}$.
We start by observing that $\frac{\varphi(r-1)}{r}\leq\frac{1}{2}$ for all primes $r$. This is clear for $r=2$, while when $r$ is odd, $\frac{\varphi(r-1)}{r}<\frac{\varphi(r-1)}{r-1}=\prod_{p\mid r-1}\big{(}1-\frac{1}{p}\big{)}\leq\frac{1}{2}$. Now for each real $\theta$ with $|\theta|\leq\frac{1}{2}$, and each positive integer $M$, we have $\log(1-\theta)=-\sum_{m\leq M}\frac{\theta^{m}}{m}+O(2^{-M})$. Thus, if we define $L_{k}$ by the equation $\prod_{r\leq r_{k}}(1-\frac{\varphi(r-1)}{r})=\exp(-L_{k})$, then
$L_{k}=\sum_{m\leq M}\frac{1}{m}\sum_{r\leq r_{k}}\left(\frac{\varphi(r-1)}{r}\right)^{m}+O(2^{-M}k).$
Here $M$ is a positive integer parameter at our disposal.
Continuing, note that $(\frac{\varphi(r-1)}{r})^{m}-(\frac{\varphi(r-1)}{r-1})^{m}\ll_{M}\frac{1}{r}$ for all primes $r$ and all positive integers $m\leq M$. Hence, for all $k\geq 2$,
$L_{k}$ $=\sum_{m\leq M}\frac{1}{m}\sum_{r\leq r_{k}}\left(\frac{\varphi(r-1)}{r-1}\right)^{m}+O_{M}\left(\sum_{r\leq r_{k}}r^{-1}\right)+O(2^{-M}k)$
$=\sum_{m\leq M}\frac{1}{m}\sum_{r\leq r_{k}}\left(\frac{\varphi(r-1)}{r-1}\right)^{m}+O_{M}(\log\log{r_{k}})+O(2^{-M}k).$ (5.2)
According to Lemma 4.4 of [19], if we set
$\sigma_{m}:=\prod_{p}\left(1-\frac{p^{m}-(p-1)^{m}}{p^{m+1}-p^{m}}\right),$ (5.3)
then
$\sum_{r\leq r_{k}}\left(\frac{\varphi(r-1)}{r-1}\right)^{m}$ $=\sigma_{m}r_{k}/\log{r_{k}}+O_{m}(r_{k}/(\log{r_{k}})^{2}).$
(In [19, Lemma 4.4], the moments of $\varphi(r-1)/(r-1)$ are estimated excluding $r=2$; including $r=2$ does not change the asymptotics.) By the prime number theorem with the de la Vallée-Poussin error bound,
$\sigma_{m}r_{k}/\log{r_{k}}+O_{m}(r_{k}/(\log{r_{k}})^{2})$ $=\sigma_{m}\pi(r_{k})+O_{m}(r_{k}/(\log{r_{k}})^{2})$
$=k\sigma_{m}+O_{m}(k/\log{k}).$
Substituting into our earlier expression (5.2) for $L_{k}$ yields
$L_{k}=k\sum_{m\leq M}\frac{\sigma_{m}}{m}+O_{M}(k/\log{k})+O(2^{-M}k).$ (5.4)
Inspecting the product definition (5.3) of $\sigma_{m}$, we see that $0<\sigma_{m}<2^{-m}$ for each positive integer $m$. (Note that the $p=2$ term in (5.3) is precisely $2^{-m}$.) It follows immediately that $\sum_{m=1}^{\infty}\frac{\sigma_{m}}{m}$ converges to a positive number $\varrho_{0}$, say. Dividing (5.4) through $k$ and sending $k$ to infinity, we find that both $\limsup_{k\to\infty}L_{k}/k$ and $\liminf_{k\to\infty}L_{k}/k$ are within $O(2^{-M})$ of $\sum_{m\leq M}\sigma_{m}/m$. Sending $M$ to infinity, we conclude that $\lim_{k\to\infty}L_{k}/k=\rho_{0}$. That is,
$L_{k}=(1+o(1))k\rho_{0},\quad\text{as}\quad k\to\infty.$
So if we define $\varrho:=\exp(\varrho_{0})$, then
$\prod_{r\leq r_{k}}\left(1-\frac{\varphi(r-1)}{r}\right)=\exp(-L_{k})=\varrho^{-(1+o(1))k},$
as desired. ∎
Put $L=\log{x}/\log\log{x}$. Let $\delta_{p}$ be defined as in (1.6), and set $M_{p}=\prod_{r\leq p}r$. Then $p_{g}=p$ precisely when $g$ belongs to one of $\delta_{p}M_{p}$ residue classes modulo $M_{p}$. Since $M_{p}\ll 3^{p}$,
$\{g:|g|\leq x:p_{g}=p\}=2\delta_{p}x+O(3^{p}).$
As $([-x,x]\setminus\mathcal{G})\ll x^{1/2}$, it follows that
$\sum_{\begin{subarray}{c}g\in\mathcal{G}\\ |g|\leq x\\ p_{g}\leq L\end{subarray}}p_{g}=\sum_{p\leq L}p\sum_{\begin{subarray}{c}g\in\mathcal{G}\\ |g|\leq x\\ p_{g}=p\end{subarray}}1$ $=\sum_{p\leq L}p\left(\Bigg{(}\sum_{\begin{subarray}{c}|g|\leq x\\ p_{g}=p\end{subarray}}1\Bigg{)}+O(x^{1/2})\right)$
$=2x\sum_{p\leq L}p\delta_{p}+O\bigg{(}\sum_{p\leq L}p(3^{p}+x^{1/2})\bigg{)}$
$=2x\sum_{p\leq L}p\delta_{p}+O(x^{1/2}L^{2}).$ (5.5)
We now extend the sum on $p$ to infinity, using Lemma 5.1 to estimate the resulting error. By (5.1),
$\delta_{p}=\frac{\varphi(p-1)}{p}\prod_{r<p}\left(1-\frac{\varphi(r-1)}{r}\right)\leq\prod_{r\leq p}\left(1-\frac{\varphi(r-1)}{r}\right)=\exp(-(1+o(1))(\log{\varrho})p/\log{p}),$
where the final estimate holds as $p\to\infty$. Consequently, if we fix $c=\frac{1}{2}\log\varrho$ (for instance), then $\delta_{p}\ll\exp(-cp/\log{p})$ for all primes $p$, and
@@
Referring back to (5.5), we deduce that
$\sum_{\begin{subarray}{c}g\in\mathcal{G}\\ |g|\leq x\\ p_{g}\leq L\end{subarray}}p_{g}=2x\sum_{p}p\delta_{p}+O(x\exp(-(\log{x})^{1+o(1)})).$ (5.6)
Next, we bound the sum of the $p_{g}$ taken over $g\in\mathcal{G}$, $|g|\leq x$, having $p_{g}>L$. If $p_{g}>L$, then $g$ belongs to one of $\xi M$ residue classes mod $M$, where
$M:=\prod_{r\leq L}r,\quad\text{and}\quad\xi:=\prod_{r\leq L}\left(1-\frac{\varphi(r-1)}{r}\right).$
The number of such $g$ with $|g|\leq x$ is $\ll\xi(x+M)\ll\xi x$, noting that $M\leq 3^{L}=x^{o(1)}$. By another application of (5.1), $\xi\leq\exp(-cL/\log{L})$. (All of this is being claimed for large enough values of $x$.) Thus,
$\{g:|g|\leq x,p_{g}>L\}\ll x\exp(-cL/\log{L}),$ (5.7)
so that by Theorem 1.2,
$\sum_{\begin{subarray}{c}g\in\mathcal{G}\\ |g|\leq x\\ p_{g}>L\end{subarray}}p_{g}$ $\leq(\max_{\begin{subarray}{c}g\in\mathcal{G}\\ |g|\leq x\end{subarray}}p_{g})\#\{g:|g|\leq x,p_{g}>L\}$
$\ll(\log{x})^{19}(x\exp(-cL/\log{L}))$
$\ll x\exp(-(\log{x})^{1+o(1)}).$
Putting together the contributions,
$\sum_{\begin{subarray}{c}g\in\mathcal{G}\\ |g|\leq x\end{subarray}}p_{g}=2x\sum_{p}p\delta_{p}+O(x\exp(-(\log{x})^{1+o(1)})).$
Corollary 1.3 follows because the function in (1.7) is $\sum_{p}p\delta_{p}+O(\exp(-(\log{x})^{1+o(1)}))$.
@@
Our main tool for this proof will be Montgomery’s “arithmetic large sieve” inequality introduced in Section 4. Using Montgomery’s sieve, Vaughan showed [19, eq. (1.3)] that for every pair of integers $M,N$ with $N>0$, we have $p_{g}\leq N^{1/2}$ for all $g\in[M+1,M+N]$ apart from $O(N^{1/2}(\log{N})^{1-\alpha})$ exceptions, where $\alpha$ is an explicit positive constant (see [19, eq. (1.4)] for its precise definition). Earlier Gallagher [6] had shown such a result with 11 in place of $1-\alpha$. The next proposition implies that $N^{1/2}$ can be replaced by a large power of $\log{N}$, if one is willing to slightly inflate the exponent $1/2$ on $N$ in the size of the exceptional set.
@@
We may assume NN is sufficiently large. We apply the arithmetic large sieve from Section 4 with $Q=N^{1/2}$, taking $\nu(p)=\varphi(p-1)$ for $p\leq Y$, and $\nu(p)=0$ for $Y<p\leq Q$. It suffices to show that with these choices of parameters, the denominator
$J=\sum_{\begin{subarray}{c}n\leq N^{1/2}\\ P^{+}(n)\leq Y\end{subarray}}\mu^{2}(n)\prod_{p\mid n}\frac{\varphi(p-1)}{p-\varphi(p-1)}$ (6.1)
@@
$J\geq N^{1/2}\exp\left(O\left(\log{N}\frac{\log\log\log{N}}{\log\log{N}}\right)\right)\cdot\exp(-u\log{u}).$ (6.2)
Let $R$ be the number of primes $p\in[\frac{1}{2}Y,Y]$ for which the smallest prime factor of $\frac{p-1}{2}$ exceeds $Y^{1/5}$. By the linear sieve and the Bombieri–Vinogradov theorem, $R\gg Y/(\log{Y})^{2}$. (This application of the linear sieve is ‘isomorphic’ to the one described at the start of [3, Chapter 8]. Here $1/5$ may be replaced by any constant smaller than $1/4$.) Let $p$ be one of these $R$ primes. Then $\frac{\varphi(p-1)}{p-1}=\frac{1}{2}\prod_{\ell\mid p-1,~\ell>2}(1-1/\ell)>\frac{1}{2}(1-y^{-1/5})^{4}>2/5$ (for instance). Hence, $\frac{\varphi(p-1)}{p}>\frac{1}{3}$, and $\frac{\varphi(p-1)}{p-\varphi(p-1)}>\frac{1}{2}$. Let $u_{0}=\lfloor\log(N^{1/2})/\log{Y}\rfloor$ (so that $u_{0}=\lfloor u\rfloor$, with $u$ as in the statement of Proposition 6.1). By considering the contribution to the right-hand side of (6.1) from products of $u_{0}$ distinct primes $p$ of the above kind, we see that $J\geq 2^{-u_{0}}\binom{R}{u_{0}}$. Now $R>Y/(\log{Y})^{3}>(\log{N})^{3/2}>u_{0}$. Since
$\binom{n}{k}=\prod_{0\leq j<k}\frac{n-j}{k-j}\geq\left(\frac{n}{k}\right)^{k}$
for each pair of integers $n,k$ with $n\geq k>0$, we conclude that
@@
The assumed bounds on $Y$ ensure that
$Y^{-1}(2(\log{Y})^{3})^{-u}=\exp\left(O\left(\log{N}\frac{\log\log\log{N}}{\log\log{N}}\right)\right).$
Our desired lower estimate (6.2) then follows by combining the last two displays. ∎
@@
Fix $K\geq 2$ with $\eta+\frac{1}{2}(1+1/K)<1$. We start by estimating the contribution of $g\in\mathcal{G}$, $|g|\leq x$, having $p_{g}\leq\log^{K}{(3x)}$.
Let $L=\log{x}/\log\log{x}$. We showed in (5.6) that (as $x\to\infty$)
@@
Furthermore (see (5.7)), the count of $g\in\mathcal{G}$, $|g|\leq x$ with $p_{g}>L$ is $O(x\exp(-cL/\log{L}))$, where $c=\frac{1}{2}\log\varrho>0$. Hence,
@@
Combining the last two displays,
$\sum_{\begin{subarray}{c}g\in\mathcal{G},~|g|\leq x\\ p_{g}\leq\log^{K}(3x)\end{subarray}}\min\{p_{g},x^{\eta}\}=\sum_{\begin{subarray}{c}g\in\mathcal{G},~|g|\leq x\\ p_{g}\leq\log^{K}(3x)\end{subarray}}p_{g}=2x\sum_{p}p\delta_{p}+O(x\exp(-(\log{x})^{1+o(1)})),$
as $x\to\infty$.
Therefore, the proof of Theorem 1.4 will be completed once it is shown that
$\sum_{\begin{subarray}{c}g\in\mathcal{G},~|g|\leq x\\ p_{g}>\log^{K}(3x)\end{subarray}}\min\{p_{g},x^{\eta}\}=o(x),$
as $x\to\infty$. For this we apply Proposition 6.1. Choose $M$ and NN with $M+1=-\lfloor x\rfloor$ and $M+N=\lfloor x\rfloor$; then $[M+1,M+N]$ is the set of all integers $g$ with $|g|\leq x$, and $N=2\lfloor x\rfloor+1<3x$. Thus, if $p_{g}>\log^{K}{(3x)}$, then $p_{g}>\log^{K}{N}$. By Proposition 6.1, the number of such $g$, $|g|\leq x$, is at most $x^{\frac{1}{2}(1+1/K)+o(1)}$. It follows that the sum appearing in the last display is bounded above by $x^{\eta}\cdot x^{\frac{1}{2}(1+1/K)+o(1)}$, which is $o(x)$ by our choice of $K$. ∎
Remark.
Our proof of Theorem 1.4 does not make essential use of the restriction $g\in\mathcal{G}$: If we average $\min\{p_{g},x^{\eta}\}$ over all integers $g$ with $|g|\leq x$, the same arguments show that the limit is again $\sum_{p}p\delta_{p}$. For this unrestricted average, $\frac{1}{2}$ is a natural boundary for $\eta$, in that even, square values of $g$ will send the average of $\min\{p_{g},x^{\frac{1}{2}+\varepsilon}\}$ to infinity for any fixed $\varepsilon>0$. One might hope to push $\eta$ past $\frac{1}{2}$ after restoring the condition that $g\in\mathcal{G}$, but it is not clear how to work the restriction of $g$ to $\mathcal{G}$ into the proof of a result like Proposition 6.1.